what is the maximum no. of phases for a two component system? Justify with example.

Grade:12th pass

1 Answers

Arun
5 years ago
To answer your question, the maximum number of phases that can be at equilibrium is when the degrees of freedom are set to 0. this is easily seen by rearranging the phase rule: P = 2 + N - F. 
The degrees of freedom have a 'negative' effect on the number of phases. By setting F = 0 and N = 2 for a two component system, the maximum number of phases is P = 2 + 2 = 4;